# Tensor network influence functionals in the continuous-time limit: connections to quantum embedding, bath discretization, and higher-order time propagation

### v1

@@ II.3 Schmidt Decomposition of Slater determinants
$\mathinner{|{\psi}\rangle}=\prod_{p=1}^{N_{\text{occ}}}\hat{c}_{p}^{{\dagger}}\mathinner{|{0}\rangle},\quad\hat{c}^{{\dagger}}_{p}=\sum_{i}C_{ip}\hat{a}^{{\dagger}}_{i}$ (16)
where $\hat{c}^{{\dagger}}_{p}$ is a creation operator of the $N_{\text{occ}}$ occupied orbitals, $\hat{a}^{{\dagger}}_{i}$ is a creation operator of orthonormal orbitals in the basis of $n$ sites, and $C_{ip}$ is the orbital coefficient matrix. Given a bipartite Hilbert space, $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, where the first $n_{A}$ orbitals belong to subsystem A and the other $n_{B}=n-n_{A}$ orbitals belong to subsystem B, the Schmidt decomposition of the Slater determinant can be obtained by diagonalizing the one-particle reduced density matrices (1-RDM), $\Gamma_{ij}=\mathinner{\langle{\psi}|}\hat{a}^{{\dagger}}_{j}\hat{a}_{i}\mathinner{|{\psi}\rangle}$, of the subsystems. Assuming $n_{A}<n_{B}$ and $N_{\text{occ}}>n_{A}$, the Schmidt decomposition can be written as,
$\mathinner{|{\psi}\rangle}=\prod_{k=1}^{n_{A}}\left(\sqrt{\nu_{k}}\hat{c}^{{\dagger}}_{A,k}+\sqrt{1-\nu_{k}}\hat{c}^{{\dagger}}_{B,k}\right)\prod_{l=n_{A}+1}^{N_{\text{occ}}}\hat{c}^{{\dagger}}_{B,l}\mathinner{|{0}\rangle}$ (17)
@@
Note that the Slater determinant with all $\nu_{k}=\frac{1}{2}$ corresponds to a maximally entangled fermionic state, $\mathinner{|{\phi}\rangle}$, where the reduced density operator of subsystem $A$ is proportional to the identity. It is possible to write the 1-RDM of $A$ from $\mathinner{|{\psi}\rangle}$ as that of a maximally entangled fermionic state $\mathinner{|{\phi}\rangle}$ after a gauge transformation within the subsystem $A$, i.e. $\hat{G}_{A}\mathinner{|{\phi}\rangle}$, where
@@
up to a normalization constant factor, assuming $10<\nu_{k}<$ 1 for all $k$. $\hat{G}_{A}$ satisfies the following
$\hat{G}_{A}\hat{c}^{{\dagger}}_{A,k}\hat{G}_{A}^{-1}=g_{k}\hat{c}^{{\dagger}}_{A,k},\>\hat{G}_{A}\hat{c}_{A,k}\hat{G}_{A}^{-1}=g_{k}^{-1}\hat{c}_{A,k}$ (19)
This gauge transformation is related to the gauge transformation introduced in Sec. II.2 because the state $\mathinner{|{\phi}\rangle}$ is ‘left-normalized’ with respect to the subsystem $A$. We will use this gauge transformation to convert Slater determinants into left-normalized forms in the next section.
@@
After a truncation to $n_{\text{ent}}$ entangled orbitals, the Slater determinant can be written as,
@@ II.4 Influence Functional tensors as Slater determinants
Figure 3: (a) A right IF state after one time-step, $\mathinner{|{\Psi^{R,1}}\rangle\!\rangle}$. It represents a partial contraction of the bath degrees of freedom between the time evolution operator, $\hat{U}^{SB}$, and the initial bath density operator, $\rho_{B}(0)$. It can be expressed as a Slater determinant by introducing a maximally entangled fermionic state, $\phi$, which has impurity and auxiliary fermionic orbitals denoted by $S$ and $A$, respectively. Coupling the maximally entangled state to the time evolution operator on the input impurity orbitals yields the Slater determinant state. (b) A right IF state after the $m$-th time-step, $\mathinner{|{\Psi^{R,m}}\rangle\!\rangle}$, ($m=3$ in the figure) can also be represented as a Slater determinant by introducing maximally entangled fermionic states at each time-step. (c) A left IF state, $\mathinner{|{\Psi^{L,m}}\rangle\!\rangle}$, ($m=N_{t}-3$ in the figure) can be constructed by contracting the time evolution operator from the top downwards, and can also be represented as a Slater determinant by inserting maximally entangled fermionic states starting from the top.
We define the right and left bipartitions of IF (right and left IF for short) $\prod_{m}\hat{U}_{m}^{SB}\cdot\mathinner{|{\rho_{B}(0)}\rangle\!\rangle}$ and $\mathinner{\langle\!\langle{\operatorname{Tr}_{B}}|}\cdot\prod_{m}\hat{U}_{m}^{SB}$, respectively, where the products follow the order in the MPS representation and here, $\cdot$ indicates the partial contraction of the intermediate bath configurations. We start with the first right IF, $\hat{U}^{SB}\cdot\mathinner{|{\rho_{B}(0)}\rangle\!\rangle}$ for simplicity. Its tensor elements are determined from $\mathinner{\langle\!\langle{s_{1}^{f},b_{1}}|}\hat{U}^{SB}(\mathinner{|{s_{1}^{i}}\rangle\!\rangle}\otimes\mathinner{|{\rho_{B}(0)}\rangle\!\rangle})$, and its external degrees of freedom are $s_{1}^{i}$, $s_{1}^{f}$, and $b_{1}$ (see Fig. 3(a)). Its occupation numbers satisfy the relationship,
@@
After applying the particle-hole transformation to the input configuration, $s_{1}^{i}\rightarrow\bar{s}_{1}^{i}$, (for example, $00\rightarrow 11$, $01\rightarrow 10$, $10\rightarrow 01$, and $11\rightarrow 00$), we obtain,
@@
where $\hat{I}^{A}$ is the identity operator on the auxiliary fermionic orbitals. Because the initial state, $\mathinner{|{\phi}\rangle\!\rangle}\otimes\mathinner{|{\rho_{B}(0)}\rangle\!\rangle}$ is given by a Slater determinant and $\hat{I}^{A}\otimes\hat{U}^{SB}$ is number-conserving, we see again that the right IF state is also a Slater determinant. The right IF state for $\prod_{m}\hat{U}_{m}^{SB}\cdot\mathinner{|{\rho_{B}(0)}\rangle\!\rangle}$ can be similarly expressed by inserting maximally entangled fermionic states at each time-step and coupling one of the modes to $\hat{U}^{SB}$, as shown schematically in Fig. 3b.
To construct the compression of the IF-MPS later, we require the bath 1-RDMs. Given the right IF state at the $m$-th time-step, $\mathinner{|{\Psi^{R,m}}\rangle\!\rangle}$, we define the bath 1-RDM of the right IF state (Fig. 4a) as,
@@
where the indices $i$ and $j$ refer to the super-Fock bath orbitals. The 1-RDM at the next time-step can be computed from the evolution of $\mathinner{|{\phi}\rangle\!\rangle}\otimes\mathinner{|{\Psi^{R,m}}\rangle\!\rangle}$ under $\hat{I}^{A}\otimes\hat{U}^{SB}$. The bath 1-RDM after the evolution can be written as,
@@
We can construct the left IF state in an analogous way, but where the state propagates in the inverse (negative) time direction using $\hat{U}^{SB{\dagger}}$ (Fig. 4c). Denoting the left IF state and its bath 1-RDM, $\mathinner{|{\Psi^{L}}\rangle\!\rangle}$ and $\Gamma^{L}$, respectively, the bath 1-RDMs at successive time-steps (from top downwards) are related by
$\Gamma^{L,m-1}_{ij}=\left[U^{SB{\dagger}}\Gamma^{L,m}U^{SB}\right]_{ij}+\left[U^{ASB{\dagger}}\Gamma^{\phi}U^{ASB}\right]_{ij},$ (27)
Note that the initial state is given by the trace vector, $\mathinner{|{\Psi^{L,N_{t}}}\rangle\!\rangle}=\mathinner{|{\operatorname{Tr}_{B}}\rangle\!\rangle}$, which is also given by a Slater determinant as in Eq. 7. With both the left and right IF state, the IF-MPS can be written as $\mathinner{\langle\!\langle{\Psi^{L,m}}|}\cdot\mathinner{|{\Psi^{R,m}}\rangle\!\rangle}$ for any $m$.
Figure 4: (a) The bath 1-RDM of the right IF state, $\Gamma^{R}$, is obtained by tracing out the impurity orbitals, i.e. by contracting the state with its complex conjugate, within the impurity space. (b) The bath 1-RDM of the right IF state after the $m$-th time-step, $\Gamma^{R,m}$ can be updated to $\Gamma^{B,m+1}$ by applying the time evolution operator and tracing out the impurity orbitals. (c) A gauge transformation, $\hat{G}$, is extracted from the bath 1-RDM of the left IF state, $\Gamma^{L}$, by converting it to a maximally entangled fermionic state, $\phi$. The gauge-transformed right IF state is obtained by applying the gauge transformation, $\hat{G}$, to the bath, and its bath 1-RDM is transformed to $\Gamma^{G}$. (d) The impurity-bath time evolution operator, $\hat{U}_{SB}$, is approximated by an effective time evolution operator by projecting the bath to the effective bath orbitals after the gauge transformation, $\mathcal{P}_{m+1}\hat{G}_{m+1}\hat{U}^{SB}\hat{G}_{m}^{-1}\mathcal{P}_{m}$. The filled and unfilled rectangles represent the core and virtual orbital spaces and these spaces are projected into fully occupied and unoccupied states, respectively.
@@ II.5 Influence Functional Matrix Product State Compression
We now revisit the MPS compression of the IF-MPS, described in Sec. II.2, but now utilizing the non-interacting nature of the bath, which allows all the steps to be expressed at the level of orbitals and single-particle quantities.
In Sec II.2, the optimal compression of the MPS required the IF-MPS to be in a canonical form. This was achieved by transforming matrices into left-normalized matrices by inserting gauge matrices into the MPS. As discussed in Sec. II.4 the partitions of the IF-MPS for a non-interacting bath are Slater determinants, and the gauge transformation (Eq. 18) to convert Slater determinants to maximally entangled fermionic pairs, which are left-normalized, was introduced in Sec. II.3.
We therefore have all the ingredients to convert the IF-MPS to canonical form. We start with the left IF state, $\mathinner{|{\Psi^{L}}\rangle\!\rangle}$, and determine the gauge transformation, $\hat{G}$, from the eigenvalues and eigenvectors of the bath 1-RDM, $\Gamma^{L}_{ij}$. We denote the eigenvalues of $\Gamma^{L}_{ij}$ as $\nu_{k}$, where $k$ indexes the eigenvalues, and the rotation matrix as $R^{L}_{ik}$, whose columns are the eigenvectors of $\Gamma^{L}$ in a single-particle basis. With this gauge transformation, we can represent the IF-MPS as, $\mathinner{\langle\!\langle{\Psi^{L}}|}\cdot\mathinner{|{\Psi^{R}}\rangle\!\rangle}=\mathinner{\langle\!\langle{\phi}|}\cdot\hat{G}\cdot\mathinner{|{\Psi^{R}}\rangle\!\rangle}$. The gauge matrix in the single-particle basis, $G$, can be written as,
@@
The gauge matrices need to be absorbed into the right IF state to further canonicalize the IF-MPS. We call the state, $\mathinner{|{\Psi^{G}}\rangle\!\rangle}=\hat{G}\cdot\mathinner{|{\Psi^{R}}\rangle\!\rangle}$, a gauge-transformed right IF state and its bath 1-RDM, $\Gamma^{G}$. The gauge transformation is a non-unitary transformation, so we always normalize the state when computing $\Gamma^{G}$.
@@
where $\Gamma^{R}$ is the bath 1-RDM of the $\mathinner{|{\Psi^{R}}\rangle\!\rangle}$. A detailed derivation of this expression is included in the Supplemental Material.
@@
Hence, the procedure for IF-MPS compression at the orbital level is as follows: (1) Diagonalize the bath 1-RDM of the gauge-transformed right IF state, $\Gamma^{G}$, and obtain its eigenvalues, $\nu_{k}^{G}$, and rotation matrix, $R^{G}$. (2) Take the eigenvectors with the $N_{\text{eff}}$ largest $\nu_{k}^{G}(1-\nu_{k}^{G})$, where $N_{\text{eff}}$ is a number of the effective bath orbitals we select, defining the truncated rotation matrix $R^{\text{eff}}$. (3) Construct the projection operator, $\mathcal{P}$, from $R^{\text{eff}}$, i.e. the eigenvectors corresponding to the effective bath orbitals. The configurations from the other bath orbitals in the projected subspace are fixed to be either fully occupied (core orbitals, $\nu_{k}^{G}\approx 1$) or unoccupied (virtual orbitals, $\nu_{k}^{G}\approx 0$).
For the Liouville time evolution of the Anderson impurity model with the initial Gaussian thermal bath, it is possible to prove that if $\nu_{k}^{G}$ is an eigenvalue of ${\Gamma}^{G}$, so is $1-\nu_{k}^{G}$ (Supplemental Material). In this case, we obtain the same number of core and virtual orbitals and choose the effective bath orbitals symmetrically by occupancy.
The tensor elements of the IF-MPS can then be computed from Eq. 15, after applying the gauge transformation and the projectors, $\mathcal{P}_{m}$, for each $m$-th time-step,
@@
with the configurations of the effective bath orbitals, $\tilde{b}_{m}$. This tensor element defines an effective time evolution operator for the embedded wavefunction defined on the impurity and effective bath orbitals. The diagrammatic representation for the effective time evolution operator is illustrated in Fig. 4d.
The above tensor elements can be efficiently computed from determinant formulae. Since the configurations in the core and virtual orbitals are fixed, we can compute the determinant of block matrices, keeping the core and virtual orbital block matrices fixed. Therefore, the computational complexity to compute determinants for all configurations is $\mathcal{O}(N_{b}^{3}+2^{2N_{\text{eff}}}N_{\text{eff}}^{3})$, where the first term corresponds to the computation of the determinant and inverse of the core and virtual orbital block matrices and the second term corresponds to the computation of the determinant of the effective bath orbital block matrices. The cost for computing the full set of MPS tensor elements at time-step $N_{t}$ is $\mathcal{O}(N_{b}^{3}N_{t}+2^{2N_{\text{eff}}}N_{\text{eff}}^{3}N_{t})$, which is linear in the number of time-steps $N_{t}$.

### v2

@@ II.3 Schmidt decomposition of Slater determinants
$\mathinner{|{\psi}\rangle}=\prod_{p=1}^{N_{\text{occ}}}\hat{c}_{p}^{{\dagger}}\mathinner{|{0}\rangle},\quad\hat{c}^{{\dagger}}_{p}=\sum_{i}C_{ip}\hat{a}^{{\dagger}}_{i},$ (16)
where $\hat{c}^{{\dagger}}_{p}$ is a creation operator of the $N_{\text{occ}}$ occupied orbitals, $\hat{a}^{{\dagger}}_{i}$ is a creation operator of orthonormal orbitals in the basis of $n$ sites, and $C_{ip}$ is the orbital coefficient matrix. Given a bipartite Hilbert space, $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, where the first $n_{A}$ orbitals belong to subsystem A and the other $n_{B}=n-n_{A}$ orbitals belong to subsystem B, the Schmidt decomposition of the Slater determinant can be obtained by diagonalizing the one-particle reduced density matrices (1-RDM), $\Gamma_{ij}=\mathinner{\langle{\psi}|}\hat{a}^{{\dagger}}_{j}\hat{a}_{i}\mathinner{|{\psi}\rangle}$, of the subsystems. Assuming $n_{A}<n_{B}$ and $N_{\text{occ}}>n_{A}$, the Schmidt decomposition can be written as
$\mathinner{|{\psi}\rangle}=\prod_{k=1}^{n_{A}}\left(\sqrt{\nu_{k}}\hat{c}^{{\dagger}}_{A,k}+\sqrt{1-\nu_{k}}\hat{c}^{{\dagger}}_{B,k}\right)\prod_{l=n_{A}+1}^{N_{\text{occ}}}\hat{c}^{{\dagger}}_{B,l}\mathinner{|{0}\rangle},$ (17)
@@
Note that the Slater determinant with all $\nu_{k}=\frac{1}{2}$ corresponds to a maximally entangled fermionic state $\mathinner{|{\phi}\rangle}$, where the reduced density operator of subsystem $A$ is proportional to the identity. It is possible to write the 1-RDM of $A$ from $\mathinner{|{\psi}\rangle}$ as that of a maximally entangled fermionic state $\mathinner{|{\phi}\rangle}$ after a gauge transformation within the subsystem $A$, i.e., $\hat{G}_{A}\mathinner{|{\phi}\rangle}$, where
@@
up to a normalization constant factor, assuming $10<\nu_{k}<$ 1 for all $k$. $\hat{G}_{A}$ satisfies the following:
$\hat{G}_{A}\hat{c}^{{\dagger}}_{A,k}\hat{G}_{A}^{-1}=g_{k}\hat{c}^{{\dagger}}_{A,k},\>\hat{G}_{A}\hat{c}_{A,k}\hat{G}_{A}^{-1}=g_{k}^{-1}\hat{c}_{A,k}.$ (19)
This gauge transformation is related to the gauge transformation introduced in Sec. II.2 because the state $\mathinner{|{\phi}\rangle}$ is “left-normalized” with respect to the subsystem $A$. We will use this gauge transformation to convert Slater determinants into left-normalized forms in the next section.
@@
After a truncation to $n_{\text{ent}}$ entangled orbitals, the Slater determinant can be written as
@@ II.4 Influence functional tensors as Slater determinants
Figure 3: (a) A right IF state after one time step $\mathinner{|{\Psi^{R,1}}\rangle\!\rangle}$. It represents a partial contraction of the bath degrees of freedom between the time-evolution operator $\hat{U}^{SB}$ and the initial bath density operator $\rho_{B}(0)$. It can be expressed as a Slater determinant by introducing a maximally entangled fermionic state, $\phi$, which has impurity and auxiliary fermionic orbitals denoted by $S$ and $A$, respectively. Coupling the maximally entangled state to the time-evolution operator on the input impurity orbitals yields the Slater determinant state. (b) A right IF state after the $m$-th time step $\mathinner{|{\Psi^{R,m}}\rangle\!\rangle}$, ($m=3$ in the figure) can also be represented as a Slater determinant by introducing maximally entangled fermionic states at each time step. (c) A left IF state $\mathinner{|{\Psi^{L,m}}\rangle\!\rangle}$ ($m=N_{t}-3$ in the figure) can be constructed by contracting the time-evolution operator from the top downwards, and can also be represented as a Slater determinant by inserting maximally entangled fermionic states starting from the top.
We define the right and left bipartitions of IF (right and left IF for short) $\prod_{m}\hat{U}_{m}^{SB}\cdot\mathinner{|{\rho_{B}(0)}\rangle\!\rangle}$ and $\mathinner{\langle\!\langle{\operatorname{Tr}_{B}}|}\cdot\prod_{m}\hat{U}_{m}^{SB}$, respectively, where the products follow the order in the MPS representation and here, $\cdot$ indicates the partial contraction of the intermediate bath configurations. We start with the first right IF, $\hat{U}^{SB}\cdot\mathinner{|{\rho_{B}(0)}\rangle\!\rangle}$ for simplicity. Its tensor elements are determined from $\mathinner{\langle\!\langle{s_{1}^{f},b_{1}}|}\hat{U}^{SB}(\mathinner{|{s_{1}^{i}}\rangle\!\rangle}\otimes\mathinner{|{\rho_{B}(0)}\rangle\!\rangle})$, and its external degrees of freedom are $s_{1}^{i}$, $s_{1}^{f}$, and $b_{1}$ (see Fig. 3(a)). Its occupation numbers satisfy the relationship
@@
After applying the particle-hole transformation to the input configuration, $s_{1}^{i}\rightarrow\bar{s}_{1}^{i}$, (for example, $00\rightarrow 11$, $01\rightarrow 10$, $10\rightarrow 01$, and $11\rightarrow 00$), we obtain
@@
where $\hat{I}^{A}$ is the identity operator on the auxiliary fermionic orbitals. Because the initial state $\mathinner{|{\phi}\rangle\!\rangle}\otimes\mathinner{|{\rho_{B}(0)}\rangle\!\rangle}$ is given by a Slater determinant and $\hat{I}^{A}\otimes\hat{U}^{SB}$ is number conserving, we see again that the right IF state is also a Slater determinant. The right IF state for $\prod_{m}\hat{U}_{m}^{SB}\cdot\mathinner{|{\rho_{B}(0)}\rangle\!\rangle}$ can be similarly expressed by inserting maximally entangled fermionic states at each time step and coupling one of the modes to $\hat{U}^{SB}$, as shown schematically in Fig. 3b.
To construct the compression of the IF-MPS later, we require the bath 1-RDMs. Given the right IF state at the $m$-th time step $\mathinner{|{\Psi^{R,m}}\rangle\!\rangle}$, we define the bath 1-RDM of the right IF state (Fig. 4a) as
@@
where the indices $i$ and $j$ refer to the super-Fock bath orbitals. The 1-RDM at the next time step can be computed from the evolution of $\mathinner{|{\phi}\rangle\!\rangle}\otimes\mathinner{|{\Psi^{R,m}}\rangle\!\rangle}$ under $\hat{I}^{A}\otimes\hat{U}^{SB}$. The bath 1-RDM after the evolution can be written as
@@
We can construct the left IF state in an analogous way, but where the state propagates in the inverse (negative) time direction using $\hat{U}^{SB{\dagger}}$ (Fig. 4c). Denoting the left IF state and its bath 1-RDM, $\mathinner{|{\Psi^{L}}\rangle\!\rangle}$ and $\Gamma^{L}$, respectively, the bath 1-RDMs at successive time steps (from top downwards) are related by
$\Gamma^{L,m-1}_{ij}=\left[U^{SB{\dagger}}\Gamma^{L,m}U^{SB}\right]_{ij}+\left[U^{ASB{\dagger}}\Gamma^{\phi}U^{ASB}\right]_{ij}.$ (27)
Note that the initial state is given by the trace vector $\mathinner{|{\Psi^{L,N_{t}}}\rangle\!\rangle}=\mathinner{|{\operatorname{Tr}_{B}}\rangle\!\rangle}$, which is also given by a Slater determinant as in Eq. 7. With both the left and right IF state, the IF-MPS can be written as $\mathinner{\langle\!\langle{\Psi^{L,m}}|}\cdot\mathinner{|{\Psi^{R,m}}\rangle\!\rangle}$ for any $m$.
Figure 4: (a) The bath 1-RDM of the right IF state, $\Gamma^{R}$, is obtained by tracing out the impurity orbitals, i.e., by contracting the state with its complex conjugate, within the impurity space. (b) The bath 1-RDM of the right IF state after the $m$-th time step, $\Gamma^{R,m}$ can be updated to $\Gamma^{B,m+1}$ by applying the time-evolution operator and tracing out the impurity orbitals. (c) A gauge transformation, $\hat{G}$, is extracted from the bath 1-RDM of the left IF state, $\Gamma^{L}$, by converting it to a maximally entangled fermionic state, $\phi$. The gauge-transformed right IF state is obtained by applying the gauge transformation, $\hat{G}$, to the bath, and its bath 1-RDM is transformed to $\Gamma^{G}$. (d) The impurity-bath time-evolution operator, $\hat{U}_{SB}$, is approximated by an effective time-evolution operator by projecting the bath to the effective bath orbitals after the gauge transformation, $\mathcal{P}_{m+1}\hat{G}_{m+1}\hat{U}^{SB}\hat{G}_{m}^{-1}\mathcal{P}_{m}$. The filled and unfilled rectangles represent the core and virtual orbital spaces and these spaces are projected into fully occupied and unoccupied states, respectively.
@@ II.5 Influence functional matrix product state compression
We now revisit the MPS compression of the IF-MPS, described in Sec. II.2, but now utilizing the noninteracting nature of the bath, which allows all the steps to be expressed at the level of orbitals and single-particle quantities.
In Sec II.2, the optimal compression of the MPS required the IF-MPS to be in a canonical form. This was achieved by transforming matrices into left-normalized matrices by inserting gauge matrices into the MPS. As discussed in Sec. II.4 the partitions of the IF-MPS for a noninteracting bath are Slater determinants, and the gauge transformation (Eq. 18) to convert Slater determinants to maximally entangled fermionic pairs, which are left-normalized, was introduced in Sec. II.3.
We therefore have all the ingredients to convert the IF-MPS to canonical form. We start with the left IF state $\mathinner{|{\Psi^{L}}\rangle\!\rangle}$ and determine the gauge transformation $\hat{G}$ from the eigenvalues and eigenvectors of the bath 1-RDM $\Gamma^{L}_{ij}$. We denote the eigenvalues of $\Gamma^{L}_{ij}$ as $\nu_{k}$, where $k$ indexes the eigenvalues, and the rotation matrix as $R^{L}_{ik}$, whose columns are the eigenvectors of $\Gamma^{L}$ in a single-particle basis. With this gauge transformation, we can represent the IF-MPS as, $\mathinner{\langle\!\langle{\Psi^{L}}|}\cdot\mathinner{|{\Psi^{R}}\rangle\!\rangle}=\mathinner{\langle\!\langle{\phi}|}\cdot\hat{G}\cdot\mathinner{|{\Psi^{R}}\rangle\!\rangle}$. The gauge matrix in the single-particle basis, $G$, can be written as
@@
The gauge matrices need to be absorbed into the right IF state to further canonicalize the IF-MPS. We call the state, $\mathinner{|{\Psi^{G}}\rangle\!\rangle}=\hat{G}\cdot\mathinner{|{\Psi^{R}}\rangle\!\rangle}$, a gauge-transformed right IF state and its bath 1-RDM, $\Gamma^{G}$. The gauge transformation is a nonunitary transformation, so we always normalize the state when computing $\Gamma^{G}$.
@@
where $\Gamma^{R}$ is the bath 1-RDM of the $\mathinner{|{\Psi^{R}}\rangle\!\rangle}$. A detailed derivation of this expression is included in the Supplemental Material (SM).
@@
Hence, the procedure for IF-MPS compression at the orbital level is as follows: (1) Diagonalize the bath 1-RDM of the gauge-transformed right IF state $\Gamma^{G}$ and obtain its eigenvalues $\nu_{k}^{G}$, and rotation matrix $R^{G}$. (2) Take the eigenvectors with the $N_{\text{eff}}$ largest $\nu_{k}^{G}(1-\nu_{k}^{G})$, where $N_{\text{eff}}$ is a number of the effective bath orbitals we select, defining the truncated rotation matrix $R^{\text{eff}}$. (3) Construct the projection operator $\mathcal{P}$ from $R^{\text{eff}}$, i.e., the eigenvectors corresponding to the effective bath orbitals. The configurations from the other bath orbitals in the projected subspace are fixed to be either fully occupied (core orbitals, $\nu_{k}^{G}\approx 1$) or unoccupied (virtual orbitals, $\nu_{k}^{G}\approx 0$).
For the Liouville time evolution of the Anderson impurity model with the initial Gaussian thermal bath, it is possible to prove that if $\nu_{k}^{G}$ is an eigenvalue of ${\Gamma}^{G}$, so is $1-\nu_{k}^{G}$ (see the SM). In this case, we obtain the same number of core and virtual orbitals and choose the effective bath orbitals symmetrically by occupancy.
The tensor elements of the IF-MPS can then be computed from Eq. 15, after applying the gauge transformation and the projectors $\mathcal{P}_{m}$ for each $m$-th time step,
@@
with the configurations of the effective bath orbitals $\tilde{b}_{m}$. This tensor element defines an effective time-evolution operator for the embedded wavefunction defined on the impurity and effective bath orbitals. The diagrammatic representation for the effective time-evolution operator is illustrated in Fig. 4d.
The above tensor elements can be efficiently computed from determinant formulas. Since the configurations in the core and virtual orbitals are fixed, we can compute the determinant of block matrices, keeping the core and virtual orbital block matrices fixed. Therefore, the computational complexity to compute determinants for all configurations is $\mathcal{O}(N_{b}^{3}+2^{2N_{\text{eff}}}N_{\text{eff}}^{3})$, where the first term corresponds to the computation of the determinant and inverse of the core and virtual orbital block matrices and the second term corresponds to the computation of the determinant of the effective bath orbital block matrices. The cost for computing the full set of MPS tensor elements at time step $N_{t}$ is $\mathcal{O}(N_{b}^{3}N_{t}+2^{2N_{\text{eff}}}N_{\text{eff}}^{3}N_{t})$, which is linear in the number of time steps $N_{t}$.